Can you help me with this equation. My son came up with the wrong answer and when i used the problem solver it was way to complicated and a different way than what my son was doing. We don't understand how the problem solver came up with his answer. I know the answer is x4 though. Can you show me the steps so i can see where he did it wrong?
5-3(4-2X)=4(x-3)-3

Answer by funmath(2933)

I don't know what problem solver you're using, but if I'm interpretting what you've typed in correctly, then:
5-3(4-2X)=4(x-3)-3 First distribute the -3 on the left and the 4 on the right.
5-3(4)-3(-2x)=4(x)+4(-3)-3
5-12+6x=4x-12-3 Combine your like terms that are on the the left and right.
(5-12)+6x=4x+(-12-3)
-7+6x=4x-15 Subtract 4x from both sides.
-7+6x-4x=-4x+4x-15
-7+2x=-15 Add 7 to both sides:
7-7+2x=-15+7
2x=-8 Divide both sides by 2:
2x/2=-8/2
x=-4
:
Check by substituting x=-4 into the original equation and see if both sides equal.
5-3(4-2(-4))=4(-4-3)-3
5-3(4+8)=4(-7)-3
5-3(12)=-28-3
5-36=-31
-31=-31 We're right!!!
Happy Calculating!!!